# Language in the Flow of Time: Time-Series-Paired Texts Weaved into a Unified Temporal Narrative

### v1

@@ Abstract
While many advances in time series models focus exclusively on numerical data, research on multimodal time series, particularly those involving contextual textual information commonly encountered in real-world scenarios, remains in its infancy. Consequently, effectively integrating the text modality remains challenging. In this work, we highlight an intuitive yet significant observation that has been overlooked by existing works: time-series-paired texts exhibit periodic properties that closely mirror those of the original time series. Building on this insight, we propose a novel framework, Texts as Time Series (TaTS), which considers the time-series-paired texts to be auxiliary variables of the time series. TaTS can be plugged into any existing numerical-only time series models and enable them to handle time series data with paired texts effectively. Through extensive experiments on both multimodal time series forecasting and imputation tasks across benchmark datasets with various existing time series models, we demonstrate that TaTS can enhance predictive performance and achieve outperformance without modifying model architectures.
Machine Learning, ICML
@@ 1 Introduction
Time series modeling plays an important role in a wide range of real-world applications, including domains of finance (Sezer et al., 2020; Jing et al., 2021; Zhou et al., 2020), healthcare (Zhang et al., 2024b; He et al., 2024), climate (Fu et al., 2024b), and energy systems (Kotzur et al., 2017; Jing et al., 2022, 2024c). While extensive research has focused on approaches that rely solely on the numerical values of time series (Qiu et al., 2025; Lin et al., ; Zhou et al., 2021; Wang et al., 2024b; Li et al., 2021), real-world scenarios often involve additional modalities that co-occur with the time series and can provide valuable complementary information (De Baets & Harvey, 2023; Rai et al., 2023; Kyei & Antwi, 2017; Jing et al., 2024a; Zheng et al., 2024b, a).
In such real-world scenarios like pandemic policymaking, economic planning, or investment strategies, human experts often incorporate additional sources of information in conjunction with time series data to forecast future trends and make well-informed decisions, as textual information can provide explanations, updates, or external factors that influence the underlying numerical patterns. However, research on effectively leveraging data from other modalities paired with time series remains in its early stages. In this work, we focus on time-series paired with texts at each timestamp, a common data format in real-world applications where textual descriptions are associated with time series at each timestamp in a parallel manner, as illustrated in Figure 3 (left). For instance, during a pandemic, numerical data such as infection rates and hospital admissions are often accompanied by textual information like government announcements and news reports (Cinelli et al., 2020). On the one hand, directly applying numerical-only models while ignoring the accompanying textual information can result in suboptimal performance as it overlooks valuable contextual information that may influence or explain the patterns in the time series. On the other hand, the current state-of-the-art approach (Liu et al., 2024a) treats the textual modality as “bag-of-texts”, disregarding the unique positional characteristics that time-series-paired texts may inherently possess. Such limitations raise a pivotal research question:
What unique attributes characterize time-series-paired texts, and how can they be systematically integrated to improve time series modeling and predictions?
In this paper, we pioneer the exploration of effectively leveraging paired texts to enrich time series analysis. We identify an intriguing phenomenon, which we term “Chronological Textual Resonance”: time-series-paired texts exhibit periodic patterns that closely reflect the temporal dynamics of their corresponding numerical time series. Notably, despite variations in words and expressions, the hidden representations of two texts associated with time series points separated by a period demonstrate high similarities, revealing a deeper alignment between textual and numerical modalities. We attribute this phenomenon to the fact that the paired texts, such as expert notes, news articles, social media posts, or event descriptions, are highly sensitive to and inherently evolve in response to the dynamics of the time series itself.
Building on these insights, we propose Texts as Time Series (TaTS), a simple yet effective framework for integrating paired texts to enhance multimodal time series modeling. As previous studies have shown that different variables in a multivariate time series exhibit similar periodicity properties (Zhang & Yan, 2023; Wang et al., 2024c; Yi et al., 2024), our discovery of Chronological Textual Resonance suggests that time-series-paired texts follow a similar pattern. This insight implies that paired texts can be considered as special auxiliary variables to augment the original time series. Motivated by this, TaTS first transforms the paired textual information into a lower-dimensional representation, then combines the original time series with the textual representations as new variables to form an augmented time series. This augmented time series is subsequently fed into existing time series models, allowing them to capture both numerical and textual temporal dynamics. By incorporating paired texts as auxiliary variables and utilizing established time series models to process them, TaTS effectively exploits the intrinsic temporal properties of the texts. Our TaTS framework offers two key benefits: (i) it effectively captures the evolving positional characteristics of texts paired with a time series by incorporating their projected representations into time series models; and (ii) it functions as a plug-in module, maintaining compatibility with existing time series models. Empirically, through extensive experiments over various benchmark datasets and multiple existing time series models, the proposed TaTS achieves state-of-the-art performance on both forecasting and imputation tasks, as shown in Figure 1. In summary, our contributions in this papers are as follows:
@@
Figure 1: Performance (Mean Square Error) of different modeling frameworks of time series with paired texts on the Environment dataset. Full results reported in Appendix C.1.
@@
  • We uncover a previously overlooked phenomenon, termed Chronological Textual Resonance (CTR), which demonstrates that time-series-paired texts exhibit periodic patterns closely aligned with their corresponding numerical time series.
@@
  • Based on this phenomenon, we propose a plug-and-play multimodal time series forecasting framework, Texts as Time Series (TaTS), which transforms text representations into auxiliary variables, seamlessly integrating them into existing time series models to capture the intrinsic temporal properties of paired texts.
@@
  • Experiments on diverse benchmark datasets and multiple existing time series models demonstrate that TaTS consistently achieves superior performance without requiring modifications to model architectures.
@@ 2 Preliminary
We use calligraphic letters (e.g., $\mathcal{A}$) for sets and bold capital letters for matrices (e.g., $\mathbf{A}$). For matrix indices, we use $\mathbf{A}[i,j]$ to denote the entry in the $i^{\textrm{th}}$ row and the $j^{\textrm{th}}$ column. For a vector $\mathbf{v}$, $v[i:j]$ represents the subvector consisting of elements from the $i^{\textrm{th}}$ to the $j^{\textrm{th}}$ position, inclusive. Additionally, $\mathbf{A}[i,:]$ returns the $i^{th}$ row in $\mathbf{A}$ and $\mathbf{A}[:i]$ returns the first $i$ rows of $\mathbf{A}$. In this paper, we focus on both time series forecasting and imputation.
Time Series Forecasting. A time series is denoted as $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\}\in\mathbb{R}^{T\times N}$, where $T$ represents the number of time steps and $N$ denotes the number of variables. $\mathbf{x}_{i}$ is the time series sequence of the $i^{\textrm{th}}$ variable. When $N>1$, the time series is referred to as a multivariate time series. Let $\mathbf{X}_{a:b}$ represent the time slice of the series from timestamp $a$ to $b$, i.e., $\mathbf{X}_{a:b}=\{\mathbf{x}_{1}[a:b],\mathbf{x}_{2}[a:b],\dots,\mathbf{x}_{N}[a:b]\}$. The task of time series forecasting is to predict the future $H$ steps, which can be formulated as:
where $\widehat{\mathbf{X}}_{T+1:T+H}$ denotes the forecasting results, $\mathcal{F}$ denotes the mapping function, and $\theta_{\text{forecast}}$ denotes the learnable parameters of $\mathcal{F}$.
Time Series Imputation. The goal of imputation is to estimate missing values in the observed time series $\mathbf{X}$, where the missing entries are denoted by a binary mask $\mathbf{M}\in\{0,1\}^{T\times N}$. Specifically, $\mathbf{M}_{t,n}=1$ indicates that $\mathbf{X}_{t,n}$ is observed, while $\mathbf{M}_{t,n}=0$ indicates that $\mathbf{X}_{t,n}$ is missing. The imputation task can be formulated as:
where $\widehat{\mathbf{X}}^{\text{Imputed}}$ represents the imputed time series, $\mathcal{G}$ denotes the imputation function, $\theta_{\text{impute}}$ denotes its learnable parameters, and $\odot$ represents the element-wise multiplication. The imputation process aims to recover the missing entries such that $\widehat{\mathbf{X}}^{\text{Imputed}}\approx\mathbf{X}$ with respect to the observed $\mathbf{X}$ values.
Extending Time Series with Paired Texts. Real-world time series often evolves alongside data from other modalities. In this work, we focus on the scenario where the time series is accompanied by paired textual data for each timestamp. Specifically, in addition to the numerical time series $\mathbf{X}\in\mathbb{R}^{T\times N}$, the dataset $\mathcal{D}=\{\mathbf{X},\mathbf{S}\}$ includes textual information $\mathbf{S}=\{s_{1},s_{2},\dots,s_{T}\}$, where each $s_{t}$ represents the text associated with timestamp $t$. Each $s_{t}$ is a string that can be tokenized into a sequence of tokens, i.e., $\text{Tokenize}(s_{t})=\{w_{t,1},w_{t,2},\dots,w_{t,L_{t}}\}$, where $L_{t}$ denotes the number of tokens in the text at time $t$. The textual data can be transformed into numerical representations using a textual encoder $\mathcal{H}_{\text{text}}$, such that:
@@ 3 Chronological Textual Resonance (CTR)
The association of time series at each timestamp may imbue time-series-paired texts with unique characteristics that can be effectively harnessed through an appropriate design. In this section, we unveil a phenomenon where paired texts exhibit periodicity closely aligned with that of the time series itself. This phenomenon is prevalent across various real-world domains, such as finance, social sciences, and transportation. To illustrate this phenomenon, we analyze three real-world time series datasets with paired texts (Liu et al., 2024a), including (i) Economy: The time series represents trade data of the United States, while the texts provide descriptions of the general economic conditions of the country. (ii) Social Good: The time series captures the unemployment rate in the United States, and the texts include detailed unemployment reports. (iii) Traffic: The time series reflects monthly travel volume trends from the U.S. Department of Transportation, with corresponding texts derived from traffic volume reports issued by the same department.
@@
For each dataset $\mathcal{D}=\{\mathbf{X},\mathbf{S}\}$, we employ the Fourier Transform (Nussbaumer, 1982; Sneddon, 1995) to analyze the frequency components of time series data, allowing us to identify its dominant periodic components, as illustrated by the blue curves in Figure 2. Furthermore, to examine the periodicity of texts, we embed each $s_{t}\in\mathbf{S}$ to obtain the text embedding $e_{t}$ at timestamp $t$. Then we compute their lag-similarity, defined as $d_{l}=\sum_{t}\mathrm{cos}(e_{t},e_{t+L})$ where $L$ is the lag and $\mathrm{cos}(\cdot,\cdot)$ represents the cosine similarity. If the text embeddings exhibit a significant periodic pattern, the lag-similarity $d_{l}$ will also fluctuate periodically as the lag $l$ increases (proof in Proposition A.1). Finally, we identify the major frequencies (those with the largest amplitudes) of the texts by applying FFT to the text lag-similarity, and mark them with red dashed lines, as shown in Figure 2. Detailed process is provided in Appendix A. We find that the major frequencies of the paired texts closely match those of the time series, indicating that the paired texts exhibit periodicity that is strongly aligned with the temporal dynamics of the time series.
@@
Figure 3: Illustration of the Texts as Time Series (TaTS) framework for capturing the temporal dynamics of paired texts. Motivated by the concept of chronological textual resonance, which suggests that paired texts exhibit behaviors similar to accompanying variables in a time series, TaTS transforms the paired texts into auxiliary variables. These variables augment the numerical sequence, forming a unified multimodal sequence that can be seamlessly integrated into any existing time series model. Compared to existing text-as-bag-of-words methods, TaTS can effectively capture the natural temporal dynamics within time-series-paired texts.
Method Economy Social Good Traffic
Uniformly Random (+)(+)( + ) 5.673 2.356 2.059 1.230 1.207 0.995
Normally Random 9.284 2.878 2.926 1.374 3.163 1.511
Exponentially Random 4.521 1.960 3.724 1.564 1.141 0.911
Using 1D Text only 1.995 1.404 1.315 0.853 0.714 0.797
Why should we leverage CTR? Parallel text provides complementary information and expert knowledge that can significantly enhance the understanding of time series data. To demonstrate the benefits of utilizing periodicity in the text modality, we present an illustrative example in the univariate forecasting task, where the goal is to use $\mathbf{X}_{1:T}=\{\mathbf{x}_{1}\}\in\mathbb{R}^{T\times 1}$ to predict the next $H$ values, $\widehat{\mathbf{X}}_{T+1:T+H}$. We first concatenate the text embeddings to form $\mathbf{E}=[e_{1};e_{2};\ldots;e_{T}]_{\text{dim=1}}\in\mathbb{R}^{d_{\text{text}}\times T}$, then replace $\mathbf{x}_{1}$ with only the first dimension of the text embeddings to leverage very partial text periodicity, $\mathbf{x}^{\prime}_{1}=(\mathbf{E}[1,:])^{\intercal}\in\mathbb{R}^{T\times 1}$. In other words, we rely solely on a single evolving dimension of the paired text features to forecast future time series values. As shown in Table 1, leveraging just the periodicity of a single text feature significantly outperforms random time series forecasting. These results highlight that even partial periodicity from the text modality contributes valuable insights for improving forecasting accuracy.
@@
Why does CTR happen? We present three key reasons for the observed alignment in periodicity between time series and their paired texts: (i) Shared External Drivers: Both the time series and their paired texts are often influenced by common external factors, such as seasonal changes, recurring events, or societal and economic cycles. These shared drivers naturally induce periodicity in both the numerical time series and the accompanying texts. (ii) Influence of Time Series on Texts: Paired texts often serve as contextual reflections of the underlying time series, adapting and evolving in response to numerical trends. For instance, news articles or government reports accompanying economic indicators are frequently updated in response to the numerical trends, leading to a periodic alignment between the two modalities. (iii) Texts Contain Additional Variables with Aligned Periodicity: Paired texts often include additional variables that are closely related to the original time series. For example, if the time series represents economic data such as GDP, the accompanying texts may reference related variables like stock market indices or inflation rates. These related variables often exhibit periodicity patterns aligned with the time series and affect the periodicity of the paired texts, but are not included in the time series data.
Algorithm 1 Texts as Time Series for forecasting task
0:  Time series with concurrent texts embeddings $\mathcal{D}=\{\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\},\mathbf{E}=\{e_{1},e_{2},\dots,e_{T}\}\}$ in the input training dataset; prediction length $H$.
1:  Initialize: Prepare training samples of sequence length $L$ and prediction length $H$ to be $\{\mathbf{X}^{(i)}=\mathbf{X}_{l_{i}+1:l_{i}+L},\mathbf{E}^{(i)}=\mathbf{X}_{l_{i}+1:l_{i}+L},\mathbf{Y}^{(i)}=\mathbf{X}_{l_{i}+L+1:l_{i}+L+H}\}_{i=1}^{n}$; time series model $\mathcal{F}(\cdot;\theta_{\text{forecast}})$; projector $\text{MLP}(\cdot;\theta_{\text{MLP}})$ for reducing dimensionality.
2:  while not converged do
3:     for training sample $\{\mathbf{X}^{(i)},\mathbf{E}^{(i)},\mathbf{Y}^{(i)}\}$ do
4:        Map $\mathbf{E}^{(i)}$ to $\mathbf{Z}^{(i)}$: $\mathbf{Z}^{(i)}[j]=\text{MLP}(\mathbf{E}^{(i)}[j];\theta_{\text{MLP}})$
5:        Compute $\mathbf{U}=[\mathbf{X}^{(i)};{(\mathbf{Z}^{(i)})}^{\intercal}]_{\text{dim=1}}$ as Eq (5).
6:        Forecast $\widehat{\mathbf{X}}^{(i)}=\mathcal{F}\left(\mathbf{U};\theta_{\text{forecast}}\right)[:N]$
7:
$\operatorname*{arg\,min}_{\Theta=\{\theta_{\text{forecast}},\theta_{\text{MLP}}\}}\mathcal{L}_{\text{forecast}}(\mathbf{X}^{(i)},\widehat{\mathbf{X}})^{(i)}$ as Eq. (7)
8:     end for
9:  end while
10:  Return: TaTS model parameters $\Theta=\{\theta_{\text{forecast}},\theta_{\text{MLP}}\}$

### v2

@@ Abstract
While many advances in time series models focus exclusively on numerical data, research on multimodal time series, particularly those involving contextual textual information commonly encountered in real-world scenarios, remains in its infancy. With recent progress in large language models and time series learning, we revisit the integration of paired texts with time series through the Platonic Representation Hypothesis [15], which posits that representations of different modalities converge to shared spaces. In this context, we identify that time-series-paired texts may naturally exhibit periodic properties that closely mirror those of the original time series. Building on this insight, we propose a novel framework, Texts as Time Series (TaTS), which considers the time-series-paired texts to be auxiliary variables of the time series. TaTS can be plugged into any existing numerical-only time series models and enable them to handle time series data with paired texts effectively. Through extensive experiments on both multimodal time series forecasting and imputation tasks across benchmark datasets with various existing time series models, we demonstrate that TaTS can enhance predictive performance without modifying model architectures. Code available at https://github.com/iDEA-iSAIL-Lab-UIUC/TaTS.
@@ 1 Introduction
Figure 1: Mean Square Error of modeling frameworks of time series with paired texts on the Environment dataset.
Time series modeling plays an important role in a wide range of real-world applications, including domains of finance [46, 17, 74], healthcare [65, 13], climate [10], and energy systems [24, 27, 18, 20]. While extensive research has focused on approaches that rely solely on the numerical values of time series [42, 33, 75, 37, 58, 53, 26], real-world scenarios often involve additional modalities that co-occur with the time series and can provide valuable complementary information [6, 44, 25, 54, 19, 72, 71].
In scenarios like pandemic policymaking, economic planning, or investment strategies, textual information can provide explanations, updates, or external factors that influence the underlying numerical patterns. However, research on effectively leveraging data from other modalities paired with time series remains in its early stages. In this work, we focus on time-series paired with texts at each timestamp, a common data format in real-world applications where textual descriptions are associated with time series at each timestamp in a parallel manner, as illustrated in Figure 3 (left). For instance, during a pandemic, infection rates are often accompanied by textual information like government announcements and news reports in real-time [5]. On the one hand, numerical-only models overlooks valuable contextual information that may influence or explain the patterns in the time series. On the other hand, the current state-of-the-art approach [35] treats the textual modality as “bag-of-texts", disregarding the unique positional characteristics that time-series-paired texts may inherently possess. Such limitations raise a pivotal question:
What unique attributes may characterize time-series-paired texts, and how can they be systematically integrated to improve time series modeling and predictions?
In this paper, motivated by the Platonic Representation Hypothesis (PRH) [15], we pioneer the exploration of effectively leveraging paired texts to enrich time series analysis. We identify an intriguing phenomenon, which we term “Chronological Textual Resonance” (CTR): depending on data quality, time-series-paired texts may exhibit periodic patterns that closely reflect the temporal dynamics of their corresponding numerical time series. Notably, despite variations in words and expressions, the hidden representations of two texts associated with time series points separated by a period may demonstrate high similarities, revealing a deeper alignment between textual and numerical modalities. We attribute this phenomenon to the fact that the paired texts are highly sensitive to and inherently evolve in response to the dynamics of the time series itself. Further, we introduce TT-Wasserstein, a new metric designed to measure the level of CTR and quantify alignment quality.
Building on these insights, we propose Texts as Time Series (TaTS), a simple yet effective framework for integrating paired texts to enhance multimodal time series modeling. Previous studies have shown that different variables in one multivariate time series exhibit similar periodicity properties [67, 54, 62], and CTR suggests that time-series-paired texts follow a similar pattern. This observation implies that paired texts can be considered as special auxiliary variables to augment the original time series. Motivated by this, TaTS first transforms the paired textual information into a lower-dimensional representation, then combines the original time series with the textual representations as new variables to form an augmented time series. This augmented time series is subsequently fed into existing time series models, allowing them to capture both numerical and textual dynamics. TaTS offers two key benefits: (i) it effectively captures the evolving positional characteristics of texts paired with a time series by incorporating their projected representations into time series models; and (ii) it functions as a plug-in module, maintaining compatibility with existing time series models. Empirically, through extensive experiments over various benchmark datasets and multiple existing time series models, the proposed TaTS achieves state-of-the-art performance on both forecasting and imputation tasks. Notably, we observe that a higher CTR level (i.e., a lower TT-Wasserstein) correlates with greater improvements compared to numerical-only modeling. In summary, our contributions are:
@@
  • We revisit multimodal time series with the PRH and uncover a previously overlooked phenomenon, termed Chronological Textual Resonance (CTR), that time-series-paired texts may exhibit periodic patterns closely aligned with their corresponding numerical time series. Further, we propose TT-Wasserstein to quantify the level of CTR and the alignment quality.
@@
  • Based on this phenomenon, we propose a plug-and-play multimodal time series forecasting framework, Texts as Time Series (TaTS), which transforms text representations into auxiliary variables, seamlessly integrating them into existing time series models.
@@
  • Experiments on diverse benchmark datasets and multiple existing time series models demonstrate TaTS’s superior performance without requiring modifications to model architectures.
@@ 2 Preliminary
We use calligraphic letters (e.g., $\mathcal{A}$) for sets and bold capital letters for matrices (e.g., $\mathbf{A}$). For matrix indices, $\mathbf{A}[i,j]$ denotes the entry in the $i^{\textrm{th}}$ row and the $j^{\textrm{th}}$ column. For a vector $\mathbf{v}$, $v[i:j]$ represents the sub-vector sliced from the $i^{\textrm{th}}$ to the $j^{\textrm{th}}$ position, inclusively. $\mathbf{A}[i,:]$ returns the $i^{th}$ row in $\mathbf{A}$ and $\mathbf{A}[:i]$ returns the first $i$ rows of $\mathbf{A}$. In this paper, we focus on both forecasting and imputation.
Time Series Forecasting. A time series is denoted as $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\}\in\mathbb{R}^{T\times N}$, where $T$ represents the number of time steps and $N$ denotes the number of variables. $\mathbf{x}_{i}$ is the time series sequence of the $i^{\textrm{th}}$ variable. When $N>1$, the time series is referred to as a multivariate time series. Let $\mathbf{X}_{a:b}$ represent the time slice of the series from timestamp $a$ to $b$, i.e., $\mathbf{X}_{a:b}=\{\mathbf{x}_{1}[a:b],\mathbf{x}_{2}[a:b],\dots,\mathbf{x}_{N}[a:b]\}$. The task of time series forecasting is to predict the future $H$ steps:
where $\mathcal{F}$ denotes the mapping function, and $\theta_{\text{forecast}}$ denotes the learnable parameters of $\mathcal{F}$.
Time Series Imputation. The goal of imputation is to estimate missing values in the time series $\mathbf{X}$, where the missing entries are denoted by a binary mask $\mathbf{M}\in\{0,1\}^{T\times N}$. Specifically, $\mathbf{M}_{t,n}=1$ indicates $\mathbf{X}_{t,n}$ being observed, $\mathbf{M}_{t,n}=0$ indicates $\mathbf{X}_{t,n}$ being missing. Formally,
where $\widehat{\mathbf{X}}^{\text{Imputed}}$ represents the imputed time series, $\mathcal{G}$ denotes the imputation function, $\theta_{\text{impute}}$ denotes its learnable parameters, and $\odot$ represents the element-wise multiplication. The imputation process aims to recover the missing entries such that $\widehat{\mathbf{X}}^{\text{Imputed}}\approx\mathbf{X}$ with respect to the actual $\mathbf{X}$ values.
Extending Time Series with Paired Texts. In addition to the numerical time series $\mathbf{X}\in\mathbb{R}^{T\times N}$, the dataset $\mathcal{D}=\{\mathbf{X},\mathbf{S}\}$ includes textual information $\mathbf{S}=\{s_{1},s_{2},\dots,s_{T}\}$, where each $s_{t}$ represents the text at timestamp $t$. Each $s_{t}$ is a string that can be tokenized into a sequence of tokens, i.e., $\text{Tokenize}(s_{t})=\{w_{t,1},w_{t,2},\dots,w_{t,L_{t}}\}$, where $L_{t}$ denotes the number of tokens in the text at time $t$. The textual data can be transformed into numerical representations using a textual encoder $\mathcal{H}_{\text{text}}$:
@@ 3 Chronological Textual Resonance (CTR)
The association of time series at each timestamp may imbue time-series-paired texts with unique characteristics that can be effectively harnessed through an appropriate design.
The Platonic Representation Hypothesis. PRH [15] posits that different modalities describing the same object converge towards a shared, latent representation. Extending this hypothesis, if time series and paired text both describe the same changing event, their representations can be seen as dynamic projections from a common underlying source, and should exhibit similar periodic properties.
@@
To illustrate this hypothesis for time series with paired texts, we analyze three real-world datasets [35], including (i) Economy: The time series represents trade data of the U.S., while the texts describe the general economic conditions of the country. (ii) Social Good: The time series captures the unemployment rate in the U.S., and the texts include detailed unemployment reports. (iii) Traffic: The time series reflects monthly travel volume trends from the U.S. Department of Transportation, with corresponding texts derived from traffic volume reports issued by the same department.
@@
For each dataset $\mathcal{D}=\{\mathbf{X},\mathbf{S}\}$, we employ the Fourier Transform [40, 47] to analyze the frequency components of time series data, allowing us to identify its dominant periodic components, as illustrated by the blue curves in Figure 2. Furthermore, to examine the periodicity of texts, we embed each $s_{t}\in\mathbf{S}$ to obtain the text embedding $e_{t}$ at timestamp $t$. Then we compute their lag-similarity, defined as $d_{l}=\sum_{t}\mathrm{cos}(e_{t},e_{t+L})$ where $L$ is the lag and $\mathrm{cos}(\cdot,\cdot)$ represents the cosine similarity. If the text embeddings exhibit a significant periodic pattern, the lag-similarity $d_{l}$ will also fluctuate periodically as the lag $l$ increases (proof in Proposition A.1). Finally, we identify the major frequencies (those with the largest amplitudes) of the texts by applying FFT to the text lag-similarity, and mark them with red dashed lines, as shown in Figure 2. Detailed process is provided in Appendix A. We find that, for these datasets, the major frequencies of the paired texts closely match those of the time series. Specifically, the paired texts also show periodicity of 12 (frequency 0.083) for monthly sampled time series, indicating that the paired texts exhibit periodicity that is strongly aligned with the temporal dynamics of the time series.
@@
Figure 3: Texts as Time Series (TaTS) framework. As paired texts may exhibit behaviors similar to accompanying variables in a time series, TaTS transforms the paired texts into auxiliary variables. These variables augment the numerical sequence, forming a unified multimodal sequence that can be seamlessly integrated into any existing time series model.
@@
Why may CTR happen? We present three key reasons for the observed alignment in periodicity between time series and their paired texts: (i) Shared External Drivers: Both the time series and their paired texts are often influenced by common external factors, such as seasonal changes, recurring events, or societal and economic cycles. These shared drivers naturally induce periodicity in both the numerical time series and the accompanying texts. (ii) Influence of Time Series on Texts: Paired texts often serve as contextual reflections of the underlying time series, adapting and evolving in response to numerical trends. For instance, news articles or government reports accompanying economic indicators are frequently updated in response to the numerical trends, leading to a periodic alignment between the two modalities. (iii) Texts Contain Additional Variables with Aligned Periodicity: Paired texts often include additional variables that are closely related to the original time series. For example, if the time series represents economic data such as GDP, the accompanying texts may reference related variables like stock market indices or inflation rates. These related variables often exhibit periodicity patterns aligned with the time series and affect the periodicity of the paired texts, but are not included in the time series data.
Quantifying CTR level. Of course we cannot expect every time series with paired texts at each timestamp to exhibit periodicity similarity. In some cases, texts aligned with timestamps may not provide meaningful or complementary information, for example, daily lottery winning numbers. To this end, we propose a new metric for CTR, TT-Wasserstein, defined as the Wasserstein distance between the normalized spectra of time series and texts. Formally, based on Wasserstein distance $W(P,Q)$ [21], for time series with paired texts dataset $\mathcal{D}=\{\mathbf{X},\mathbf{S}\}$, after computing normalized frequencies and amplitudes $\tilde{f}_{\text{texts}},\tilde{f}_{\text{ts}},\tilde{a}_{\text{texts}},\tilde{a}_{\text{ts}}$,
$\begin{split}\text{TT-Wasserstein}(\mathcal{D})&=W(P_{\text{texts}},P_{\text{ts}})=\inf_{\gamma\in\Pi(P_{\text{text}},P_{\text{ts}})}\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}\cdot\left|\tilde{f}_{\text{texts}}^{(i)}-\tilde{f}_{\text{ts}}^{(j)}\right|\\ \text{Subject to}&\quad\gamma_{ij}\geq 0,\quad\sum_{j=1}^{m}\gamma_{ij}=\tilde{a}_{\text{texts}}^{(i)},\quad\sum_{i=1}^{n}\gamma_{ij}=\tilde{a}_{\text{ts}}^{(j)}\end{split}$ (4)
[…]
Table 1: TT-Wasserstein measure of Time-MMD [35] datasets and their random shuffle.
Dataset Monthly Sampled Weekly Sampled Daily Sampled
Agriculture Climate Economy Security Social Good Traffic Energy Health Environment
Original 0.026 0.025 0.022 0.049 0.027 0.035 0.307 0.233 0.302
TS shuffled 0.088 0.032 0.098 0.054 0.069 0.102 0.320 0.268 0.358
Text shuffled 0.106 0.037 0.099 0.053 0.072 0.104 0.312 0.277 0.364